# Computing High-dimensional Confidence Sets for Arbitrary Distributions

### v1

@@ 4 Finding Confidence Sets $(\Gamma=1+d^{-1/3+o(1)})$-Competitive with Euclidean Balls
Algorithm Dense_Ellipsoid Input: a set of point $Y\in\mathbb{R}^{d}$, target fraction $\delta\in(0,1)$, and slack $\gamma\in(0,1)$ Output: an ellipsoid $\widehat{E}\subset\mathbb{R}^{d}$ 1. Create a list of coarse balls $\mathfrak{B}=\left\{B(y_{1},\|y_{1}-y_{2}\|)~{}|~{}y_{1},y_{2}\in Y\right\}$ and remove from $\mathfrak{B}$ all balls that contains less than $\delta n$ points in $Y$. 2. Create an ellipsoid $\widehat{E}_{i}$ for each coarse ball $B^{\prime}_{i}$ in the list $\mathfrak{B}$ as follows: (a) Set $\widehat{\tau}=d^{1/6}$ and $R^{\prime}_{i}$ be the radius of $B^{\prime}_{i}$. (b) Compute the eigenvalues $\lambda_{1},\dots,\lambda_{d}$ and eigenvectors $v_{1},\dots,v_{d}$ of the covariance $\Sigma_{Y^{\prime}_{i}}$ of points $Y^{\prime}_{i}=Y\cap B^{\prime}_{i}$. (c) Create an ellipsoidal shape $\widehat{M}_{i}=\sum_{j=1}^{d}a_{j}^{2}v_{j}v_{j}^{\top}$, where $a_{j}^{2}=d$ if $\lambda_{j}\geq\widehat{\tau}^{2}(R^{\prime}_{i})^{2}/d$ and $a_{i}^{2}=1$ if $\lambda_{j}<\widehat{\tau}^{2}(R^{\prime}_{i})^{2}/d$. (d) Compute the transformed points $Z_{i}=\widehat{M}_{i}^{-1/2}Y$. Find a ball $\widehat{B}_{i}$ using subroutine from Lemma 3.6 for $q=0$, $\lambda_{(q+1)}(\Sigma_{Y^{\prime}})\leq\sigma^{2}$, $\sigma^{2}=\widehat{\tau}^{2}(R^{\prime})^{2}/d$, slack factor $\gamma$. (e) Set $\widehat{E}_{i}=\widehat{M}_{i}^{1/2}\widehat{B}$. 3. Return the ellipsoid $\widehat{E}$ with the smallest volume in the list $\{\widehat{E}_{i}\}$.
@@
Algorithm Dense_Ellipsoid chooses an ellipsoidal shape $\widehat{M}$ to have the same eigenvectors as $\Sigma_{Y^{\prime}}$, with eigenvalues $a_{i}^{2}$ of $\widehat{M}$ satisfying $a_{i}^{2}=d$ if $\lambda_{i}\geq\widehat{\tau}^{2}(R^{\prime})^{2}/d$ and $a_{i}^{2}=1$ if $\lambda_{i}<\widehat{\tau}^{2}(R^{\prime})^{2}/d$. $\widehat{M}$ has the following nice properties.
@@
We apply the procedure from Lemma 3.6 to the minimum volume ball $B_{T}^{\prime}=B(c_{T}^{\prime},R_{T}^{\prime})$ containing $T^{\prime}$, where for $q=0$, $\lambda_{(q+1)}(\Sigma_{Y^{\prime}})\leq\sigma^{2}$, where $\sigma^{2}=\widehat{\tau}^{2}(R^{\prime})^{2}/d$ (due to (13)), and slack factor $\gamma$. Let $B_{T}^{\star}=B(c_{T}^{\star},R_{T}^{\star})$ be the minimum volume ball containing $T^{\star}$. Lemma 3.6 allows us to find a ball $\widehat{B}_{T}=B(\widehat{c}_{T},\widehat{R}_{T})$ that contains a subset $\widehat{T}\subseteq T^{\prime}$, where $|\widehat{T}|\geq(1-\gamma)|T^{\star}|$, and
$\widehat{R}_{T}\leq R_{T}^{\star}+2\sigma\left(\frac{2}{\gamma}+\sqrt{2(1-\delta)/\delta}\right).$ (15)
(Lemma 3.6 has an additional parameter also named $\tau$, that is different from $\widehat{\tau}$, but this parameter does not apply here since we set $q=0$ and avoid the grid search element of the procedure entirely.) Finally, we transform $\widehat{B}_{T}$ back to the transformed space to get the candidate ellipsoid $\widehat{E}$,
@@
$\leq\exp\left(\frac{d\ln d}{\widehat{\tau}^{2}}\right)\cdot c_{d}\left(R^{\star}_{T}+2\sigma\left(\frac{2}{\gamma}+\sqrt{2(1-\delta)/\delta}\right)\right)^{d}$ by (15)
$\leq\exp\left(\frac{d\ln d}{\widehat{\tau}^{2}}\right)\cdot c_{d}(R^{\star}_{T})^{d}\cdot\left(1+2\frac{\sigma}{R^{\star}_{T}}\left(\frac{2}{\gamma}+\sqrt{2(1-\delta)/\delta}\right)\right)^{d}$
$\leq\exp\left(\frac{d\ln d}{\widehat{\tau}^{2}}\right)\cdot\exp\left(2\frac{d\sigma}{R^{\star}_{T}}\left(\frac{2}{\gamma}+\sqrt{2(1-\delta)/\delta}\right)\right)\cdot\mathrm{vol}(B^{\star}_{T})$
$\leq\exp\left(\frac{d\ln d}{\widehat{\tau}^{2}}\right)\cdot\exp\left(2\frac{d\widehat{\tau}R^{\prime}}{R^{\star}_{T}\sqrt{d}}\left(\frac{2}{\gamma}+\sqrt{2(1-\delta)/\delta}\right)\right)\cdot\mathrm{vol}(B^{\star}_{T})$ since $\text{since }\sigma^{2}=\frac{\widehat{\tau}^{2}(R^{\prime})^{2}}{d}$
$\leq\exp\left(\frac{d\ln d}{\widehat{\tau}^{2}}+2\sqrt{d}\widehat{\tau}\left(\frac{2}{\gamma}+\sqrt{2(1-\delta)/\delta}\right)\right)\cdot\mathrm{vol}(B^{\star})$ by (11)
$\leq\exp\left(d^{2/3}\left(\ln d+2\left(\frac{2}{\gamma}+\sqrt{2(1-\delta)/\delta}\right)\right)\right)\cdot\mathrm{vol}(B^{\star})$ since $\text{since }\widehat{\tau}=d^{1/6}$
$\leq\exp\left(d^{2/3+o(1)}\cdot\left(\frac{1}{\delta}+\frac{1}{\sqrt{\gamma\delta}}\right)\right)\cdot\mathrm{vol}(B^{\star}).$
@@ 5 Greedy Algorithm and Unions of Sets
$\mathrm{vol}(\widehat{C})\leq\exp\left(O_{\delta}\left(d^{2/3+o(1)}\right)\right)\cdot O\left(\frac{1}{\gamma}\cdot\log(k/\gamma)\right)\cdot\mathrm{vol}(C^{\star}),$
@@
Let $\delta\in(0,1),\gamma\in(0,1),k\in\mathbb{N}$ be any constants. There is a polynomial time algorithm that for target coverage $\delta\in(0,1)$ and coverage slack $\gamma\in(0,1)$ when given $n=\Omega(kd^{2}/\gamma^{2})$ samples drawn i.i.d. from an arbitrary distribution $\mathcal{D}$, finds with high probability a set $S\subset\mathbb{R}^{d}$ that is $\Gamma=\left(1+O_{\gamma,\delta}\big{(}d^{-1/3+o(1)}\big{)}\right)$ competitive; more precisely, it satisfies $\mathbf{P}_{y\sim\mathcal{D}}\left[y\in S\right]\geq\delta$, and
$\operatorname*{\mathrm{vol}}(S)^{1/d}\leq\mathrm{vol}(C^{\star}_{k})^{1/d}\left(1+O_{k,\delta}\left(d^{-1/3+o(1)}\right)\right)\cdot\left(\frac{O(\log(k/\gamma))}{\gamma}\right)^{1/d}$
@@ 6.2 Computational Intractability even with Slack in Coverage
Conjecture 6.2 (SSE hypothesis of Raghavendra and Steurer (2010)).
For any $\eta>0$, there is a constant $\tau\in(0,1)$ such that there is no polynomial time algorithm to distinguish between the following two cases given a graph $G=(V,E)$ on $n$ vertices with degree $D$:
@@
  • NO: Any set $S\subseteq V$ with $|S|\leq 2\tau n$ has most of the edges incident on it going outside i.e., $|E(S,V\setminus S)|\geq(1-\eta)|S|D$
@@
Soundness argument (NO case): Suppose there exists a subset of points corresponding to vertices $S^{\prime}\subset V$ with $|S^{\prime}|=k^{\prime}$ for $3\gamma k/4\leq k^{\prime}\leq k$ that are contained in a ball of radius $R^{\prime}$ around a point $c^{\prime}\in\mathbb{R}^{n}$. Note that $|S^{\prime}|=k^{\prime}\geq 3\gamma k/4=\gamma\delta n$. We have

### v2

@@ 4 Finding Confidence Sets $(\Gamma=1+d^{-1/2+o(1)})$-Competitive with Euclidean Balls
Algorithm Dense_Ellipsoid Input: a set of point $Y\in\mathbb{R}^{d}$, target fraction $\delta\in(0,1)$, and slack $\gamma\in(0,1)$ Output: an ellipsoid $\widehat{E}\subset\mathbb{R}^{d}$ 1. Create a list of coarse balls $\mathfrak{B}=\left\{B(y_{1},\|y_{1}-y_{2}\|)~{}|~{}y_{1},y_{2}\in Y\right\}$ and remove from $\mathfrak{B}$ all balls that contains less than $\delta n$ points in $Y$. 2. Create an ellipsoid $\widehat{E}_{i}$ for each coarse ball $B^{\prime}_{i}$ in the list $\mathfrak{B}$ as follows: (a) Set $\widehat{\tau}=d^{1/4}$ and $R^{\prime}_{i}$ be the radius of $B^{\prime}_{i}$. (b) Compute the eigenvalues $\lambda_{1},\dots,\lambda_{d}$ and eigenvectors $v_{1},\dots,v_{d}$ of the covariance $\Sigma_{Y^{\prime}_{i}}$ of points $Y^{\prime}_{i}=Y\cap B^{\prime}_{i}$. (c) Create an ellipsoidal shape $\widehat{M}_{i}=\sum_{j=1}^{d}a_{j}^{2}v_{j}v_{j}^{\top}$, where $a_{j}^{2}=d$ if $\lambda_{j}\geq\widehat{\tau}^{2}(R^{\prime}_{i})^{2}/d$ and $a_{i}^{2}=1$ if $\lambda_{j}<\widehat{\tau}^{2}(R^{\prime}_{i})^{2}/d$. (d) Compute the transformed points $Z_{i}=\widehat{M}_{i}^{-1/2}Y$. Find a ball $\widehat{B}_{i}$ using subroutine from Lemma 3.6 for $q=0$, $\lambda_{(q+1)}(\Sigma_{Y^{\prime}})\leq\sigma^{2}$, $\sigma^{2}=\widehat{\tau}^{2}(R^{\prime})^{2}/d$, slack factor $\gamma$. (e) Set $\widehat{E}_{i}=\widehat{M}_{i}^{1/2}\widehat{B}$. 3. Return the ellipsoid $\widehat{E}$ with the smallest volume in the list $\{\widehat{E}_{i}\}$.
@@
For a parameter $\widehat{\tau}$ that we will set later, Algorithm Dense_Ellipsoid chooses an ellipsoidal shape $\widehat{M}$ to have the same eigenvectors as $\Sigma_{Y^{\prime}}$, with eigenvalues $a_{i}^{2}$ of $\widehat{M}$ satisfying $a_{i}^{2}=d$ if $\lambda_{i}\geq\widehat{\tau}^{2}(R^{\prime})^{2}/d$ and $a_{i}^{2}=1$ if $\lambda_{i}<\widehat{\tau}^{2}(R^{\prime})^{2}/d$. $\widehat{M}$ has the following nice properties.
@@
We apply the procedure from Lemma 3.6 with $q=0$ to the minimum volume ball $B_{T}^{\prime}=B(c_{T}^{\prime},R_{T}^{\prime})$ containing $T^{\prime}$. We have that $\lambda_{\mathrm{max}}(\Sigma_{Y^{\prime}})\leq\sigma^{2}$, where $\sigma^{2}=\widehat{\tau}^{2}(R^{\prime})^{2}/d$ (due to (13)), and slack factor $\gamma$. Let $B_{T}^{\star}=B(c_{T}^{\star},R_{T}^{\star})$ be the minimum volume ball containing $T^{\star}$. Lemma 3.6 allows us to find a ball $\widehat{B}_{T}=B(\widehat{c}_{T},\widehat{R}_{T})$ that contains a subset $\widehat{T}\subseteq T^{\prime}$, where $|\widehat{T}|\geq(1-\gamma)|T^{\star}|$, and
$\widehat{R}_{T}\leq\sqrt{(R_{T}^{\star})^{2}+\frac{\sigma^{2}}{\gamma}\left(1+\frac{2(1-\delta)}{\delta}\right)}.$ (15)
(Lemma 3.6 has an additional parameter also named $\tau$, that is different from $\widehat{\tau}$, but this parameter does not apply here since we set $q=0$ and avoid the grid search element of the procedure entirely.)
Finally, we transform $\widehat{B}_{T}$ back to the transformed space to get the candidate ellipsoid $\widehat{E}$,
@@
$\leq\exp\left(\frac{d\ln d}{\widehat{\tau}^{2}}\right)\cdot c_{d}\left((R^{\star}_{T})^{2}+\sigma^{2}\gamma\left(1+\frac{2(1-\delta)}{\delta}\right)\right)^{d/2}$ by (15)
[…]
$\leq\exp\left(\frac{d\ln d}{\widehat{\tau}^{2}}\right)\cdot c_{d}(R^{\star})^{d}\left(1+\frac{\sigma^{2}}{\gamma(R^{\star})^{2}}\left(1+\frac{2(1-\delta)}{\delta}\right)\right)^{d/2}$
$\leq\exp\left(\frac{d\ln d}{\widehat{\tau}^{2}}\right)\cdot\exp\left(\frac{d}{2}\cdot\frac{\sigma^{2}}{\gamma(R^{\star})^{2}}\left(1+\frac{2(1-\delta)}{\delta}\right)\right)\cdot\mathrm{vol}(B^{\star})$
$\leq\exp\left(\frac{d\ln d}{\widehat{\tau}^{2}}\right)\cdot\exp\left(\frac{\widehat{\tau}^{2}(R^{\prime})^{2}}{2\gamma(R^{\star})^{2}}\left(1+\frac{2(1-\delta)}{\delta}\right)\right)\cdot\mathrm{vol}(B^{\star})$ since $\text{since }\sigma^{2}=\frac{\widehat{\tau}^{2}(R^{\prime})^{2}}{d}$
$\leq\exp\left(\frac{d\ln d}{\widehat{\tau}^{2}}+\widehat{\tau}^{2}\cdot\frac{2}{\gamma}\left(1+\frac{2(1-\delta)}{\delta}\right)\right)\cdot\mathrm{vol}(B^{\star})$ since $\text{since }R^{\prime}\leq 2R^{\star}$
$\leq\exp\left(d^{1/2}\left(\ln d+\frac{2}{\gamma}\left(1+\frac{2(1-\delta)}{\delta}\right)\right)\right)\cdot\mathrm{vol}(B^{\star})$ setting $\text{setting }\widehat{\tau}^{2}=d^{1/2}$
$\leq\exp\left(d^{1/2+o(1)}(\gamma\delta)^{-1}\right)\cdot\mathrm{vol}(B^{\star}).$
@@ 5 Greedy Algorithm and Unions of Sets
$\mathrm{vol}(\widehat{C})\leq\exp\left(O_{\delta}\left(d^{1/2+o(1)}\right)\right)\cdot O\left(\frac{1}{\gamma}\cdot\log(k/\gamma)\right)\cdot\mathrm{vol}(C^{\star}),$
@@
Let $\delta\in(0,1),\gamma\in(0,1),k\in\mathbb{N}$ be any constants. There is a polynomial time algorithm that for target coverage $\delta\in(0,1)$ and coverage slack $\gamma\in(0,1)$ when given $n=\Omega(kd^{2}/\gamma^{2})$ samples drawn i.i.d. from an arbitrary distribution $\mathcal{D}$, finds with high probability a set $S\subset\mathbb{R}^{d}$ that is a union of balls, and is $\Gamma=\left(1+O_{\gamma,\delta}\big{(}\log\log d/\log d\big{)}\right)$ competitive; more precisely, it satisfies $\mathbf{P}_{y\sim\mathcal{D}}\left[y\in S\right]\geq\delta$, and
$\operatorname*{\mathrm{vol}}(S)^{1/d}\leq\mathrm{vol}(C^{\star}_{k})^{1/d}\Big{(}1+O_{k,\delta}\left(\log\log d/\log d\right)\Big{)}\cdot\left(\frac{O(\log(k/\gamma))}{\gamma}\right)^{1/d}$
@@ 6.1 NP-hardness of Proper Learning
We remark that the above hardness result rules out polynomial time proper learning algorithms that work for arbitrary distributions— this is called “the distribution-free setting” in PAC learning (as opposed to the distribution-specific setting). More formally, one can consider the following reduction from the worst-case problem of finding the minimum volume ball containing a $\delta$ fraction of an adversarially chosen point set $Y\subset\mathbb{R}^{d}$, to our problem. We construct a distribution $D_{Y}$ that is the uniform distribution over $Y$. By the construction of $Y$, finding a set that contains probability mass $\delta$ over $D_{Y}$ corresponds exactly to finding a set that contains at least a $\delta$ fraction of $Y$. We can simulate running our algorithm on $D_{Y}$ by sampling from $D_{Y}$ and providing these samples to the algorithm. Thus our problem is only harder than the worst-case problem, because it only gets sample access to $D_{Y}$, rather than access to $Y$ itself.
@@ 6.2 Computational Intractability even with Slack in Coverage
Conjecture 6.2 (SSE hypothesis of Raghavendra and Steurer (2010), see e.g., Theorem IV.5 of Raghavendra et al. (2012)).
For any constant $\eta\in(0,1/2)$, there is a constant $\tau\in(0,1)$ such that there is no polynomial time algorithm to distinguish between the following two cases given a graph $G=(V,E)$ on $n$ vertices with degree $D$:
@@
  • NO: Any set $S\subseteq V$ with $\eta\tau n\leq|S|\leq 2\tau n$ has most of the edges incident on it going outside i.e., $|E(S,V\setminus S)|\geq(1-\eta)|S|D$.
@@
Soundness argument (NO case): Suppose there exists a subset of points corresponding to vertices $S^{\prime}\subset V$ with $|S^{\prime}|=k^{\prime}$ for $3\gamma k/4\leq k^{\prime}\leq k$ that are contained in a ball of radius $R^{\prime}$ around a point $c^{\prime}\in\mathbb{R}^{n}$. Note that $|S^{\prime}|\geq\eta\tau n$ and $|S^{\prime}|=k^{\prime}\geq 3\gamma k/4=\gamma\delta n$. We have